OpenStudy (anonymous):

How do I solve an indefinite integral such as (x+4)/((x^2+8x-7)^2 ? I'm not sure how to handle the fraction...

OpenStudy (mimi_x3):

Let u = x^2 + 8x - 7 du/dx = 2x + 8 dx = du/2x +8 integral ((x+4)/ u^2) * du/ (2x +8) -> integral ((x+4)/ u^2) * du/ 2(x+4) = 1/2 integral 1/u^2 integrate it you will get -1/2u + C Then sub u - x^2 + 8x - 7 back in.

OpenStudy (mimi_x3):

u = x^2 + 8x - 7*